# Generalization Error of the Tilted Empirical Risk

### v1

@@ Appendix A Other Related Works
Appendix B Technical Tools
@@
Lemma B.4 (McDiarmid’s inequality (McDiarmid, 1989)).
@@
Lemma B.5 (Hoeffding Inequality, Boucheron et al., 2013).
@@
Lemma B.6 (Bernstein’s Inequality, Boucheron et al., 2013).
@@
Lemma B.7.
@@
Lemma B.8 (Uniform bound (Mohri et al., 2018)).
@@
Lemma B.9 (Talagrand’s contraction lemma (Shalev-Shwartz and Ben-David, 2014)).
@@
Lemma B.10 (Massart’s lemma (Massart, 2000)).
@@
Lemma B.11 (Lemma 1 in (Xu and Raginsky, 2017)).
@@
Lemma B.12 (Coupling Lemma).
@@
Lemma B.13 (Kantorovich-Rubenstein duality of total variation distance, see  (Polyanskiy and Wu, 2022)).
@@
Appendix C Proofs and Details of Section 3
C.1 Uniform bound: details for bounded loss
@@
A similar approach can be applied for $\gamma>0$ by using Lemma B.3 and the final result holds. ∎
@@
C.1.1 Another Approach for Uniform Bounds
@@
Using Lemma B.3 and Jensen’sinequality, we have for $\gamma<0$,
@@
Next we use McDiarmid’s inequality, Lemma B.4, to derive an upper bound on the absolute value of the term $I_{3}$, in the following proposition.
Proposition C.2.
@@
in McDiarmid’s inequality, Lemma B.4. Note that for the term $\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(-\gamma M)}{n}\Bigg{)}$, we need that $\frac{1-\exp(-\gamma M)}{n}>-1$, and therefore we should have $|\gamma|<\frac{\log(n+1)}{M}$; otherwise $\tilde{c}=\infty$.
@@
With Propositions  C.23.2, we derive upper and lower bounds on the tilted generalization error.
Theorem C.3.
@@
Combining the results in Proposition C.1, Proposition C.2 and Proposition 3.2, we can derive the following upper bound on generalization error,
@@
Remark C.4 (The influence of $\gamma$).
As $\gamma\rightarrow 0$, the upper bound in Theorem C.3 on the tilted generalization error converges to the upper bound on the generalization error under the ERM algorithm obtained by Shalev-Shwartz and Ben-David (2014), (16). In particular, $c=c(\gamma)\rightarrow\frac{M}{n}$ and the first term in (50) vanishes. Therefore, the upper bound converges to a uniform bound on the linear empirical risk.
Corollary C.5.
Under the same assumptions in Corollary C.5, the upper bound on the tilted generalization error in Theorem C.3 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
@@
Theorem C.6 (Uniform Lower Bound).
Under the assumptions of Theorem C.3 for a fixed $h\in\mathcal{H}$ we have
@@
Remark C.7.
For $\frac{\log(n)}{2M}<|\gamma|<\frac{\log(n+1)}{M}$, the lower bound in Theorem C.6 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error, (53), for a positive $\gamma$, due to the term $\frac{\exp(-2\gamma M)}{n}-1$ in Theorem C.6.
Remark C.8.
The terms $\frac{(1/n-\exp(-2\gamma M))}{\gamma}$ in Theorem C.3 can cause the upper bounds in Theorem C.3 to be tighter for $\gamma>0$ than for $\gamma<0$. Furthermore, in comparison with the uniform upper bound on the linear generalization error, (16), we obtain a tilted generalization error upper bound, (50), which can be tighter, for sufficiently large value of samples $n$ and small values of $\gamma$.
C.2 Information-theoretic bounds: details for bounded loss
@@
The proof follows directly from applying Lemma B.12 to the $\log(.)$ function and then applying Lemma B.11. ∎
@@
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\leq\frac{\exp(-\gamma M)}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\frac{-\exp(-\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(-\gamma M)}{2\gamma}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\gamma\exp(-\gamma M)}{2}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}\end{split}$ (56)
@@
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\geq\frac{1}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\geq\frac{-\exp(\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(\gamma M)}{2\gamma}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\gamma\exp(-\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (57)
@@
C.2.1 Another Approach for an Information-theoretic Bound
@@
Using Proposition C.9 and Lemma B.3, we can derive the following upper bound on the expected generalization error.
Theorem C.10.
@@
Using Proposition C.9, we have
@@
Appendix D Proofs and Details of Section 4
D.1 Uniform bounds: details for unbounded loss
@@
Given Assumption 4.1 and Assumption 4.2, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$, $\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\exp(-\gamma R_{T}^{u})\sqrt{\frac{M_{2}^{u}\log(2/\delta)}{n}}-\frac{4\exp(-\gamma R_{T}^{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}M_{2}^{u}.\end{split}$ (66)
@@
Using Bernstein’s inequality, Lemma B.6, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$ and considering $10<X_{i}<$ 1, we have
@@
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\exp(-\gamma R_{T}^{u})\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma R_{T}^{u})\frac{4\log(2/\delta)}{3n}.\end{split}$
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\leq\frac{\exp(-\gamma R_{T}^{u})}{|\gamma|}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma R_{T}^{u})\frac{4\log(2/\delta)}{3n|\gamma|}.\end{split}$
Using Lemma B.2, completes the proof. ∎
@@
Given Assumption 4.1 and Assumption 4.2, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}M_{2}^{u}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma R_{T}^{u})}$ and $\gamma<0$, for any fixed $h\in\mathcal{H}$, the following lower bound on the tilted generalization error holds with probability at least $(1-\delta)$; $\mathrm{gen}_{\gamma}(h,S)\geq-\frac{2\exp(-\gamma R_{T}^{u})}{(1-\zeta)}\sqrt{\frac{M_{2}^{u}\big{(}\log(2/\delta)\big{)}}{n}}+\frac{4\exp(-\gamma R_{T}^{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.$ (68)
@@
First, we apply Lemma B.7 to yield $\mathrm{R}(h,\mu)-\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])\geq 0$. Next we focus on the second line of this display. Bernstein’s inequality, Lemma B.6, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$, so that $10<X_{i}<$ 1, gives that with probability at least $(1-\delta)$,

### v2

@@ Appendix B Other Related Works
Appendix C Technical Tools
@@
Lemma C.4.
Suppose that $0<a<X<b<\infty$. Then the following inequality holds,
$\begin{split}\frac{-\mathrm{Var}_{P_{X}}(X)\exp(b)}{2}\leq\exp(E[X])-E[\exp(X)]\leq\frac{-\mathrm{Var}_{P_{X}}(X)\exp(a)}{2},\end{split}$
@@
Lemma C.5 (McDiarmid’s inequality (McDiarmid, 1989)).
@@
Lemma C.6 (Hoeffding Inequality, Boucheron et al., 2013).
@@
Lemma C.7 (Bernstein’s Inequality, Boucheron et al., 2013).
@@
Lemma C.8.
For a positive random variable, $Z>0$, suppose $\mathbb{E}[Z^{2}]\leq\eta$ Then, the following inequality holds,
$\mathbb{E}[Z]\leq\eta^{1/2}.$
Lemma C.9.
@@
Lemma C.10.
Suppose that $0<a<x<b$ and $f(x)$ is an increasing and concave function. Then the following holds,
$f^{\prime}(b)(b-a)\leq f(b)-f(a)\leq f^{\prime}(a)(b-a).$ (36)
Lemma C.11 (Uniform bound (Mohri et al., 2018)).
@@
Lemma C.12 (Talagrand’s contraction lemma (Shalev-Shwartz and Ben-David, 2014)).
@@
Lemma C.13 (Massart’s lemma (Massart, 2000)).
@@
Lemma C.14 (Lemma 1 in (Xu and Raginsky, 2017)).
@@
Lemma C.15 (Coupling Lemma).
@@
Lemma C.16 (Kantorovich-Rubenstein duality of total variation distance, see  (Polyanskiy and Wu, 2022)).
@@
Appendix D Proofs and Details of Section 3
D.1 Uniform bound: details for bounded loss
@@
A similar approach can be applied for $\gamma>0$ by using Lemma C.3 and the final result holds. ∎
@@
D.1.1 Another Approach for Uniform Bounds
@@
Using Lemma C.3 and Jensen’sinequality, we have for $\gamma<0$,
@@
Next we use McDiarmid’s inequality, Lemma C.5, to derive an upper bound on the absolute value of the term $I_{3}$, in the following proposition.
Proposition D.2.
@@
in McDiarmid’s inequality, Lemma C.5. Note that for the term $\frac{1}{\gamma}\log\Bigg{(}1+\frac{1-\exp(-\gamma M)}{n}\Bigg{)}$, we need that $\frac{1-\exp(-\gamma M)}{n}>-1$, and therefore we should have $|\gamma|<\frac{\log(n+1)}{M}$; otherwise $\tilde{c}=\infty$.
@@
With Propositions  D.23.2, we derive upper and lower bounds on the tilted generalization error.
Theorem D.3.
@@
Combining the results in Proposition D.1, Proposition D.2 and Proposition 3.2, we can derive the following upper bound on generalization error,
@@
Remark D.4 (The influence of $\gamma$).
As $\gamma\rightarrow 0$, the upper bound in Theorem D.3 on the tilted generalization error converges to the upper bound on the generalization error under the ERM algorithm obtained by Shalev-Shwartz and Ben-David (2014), (15). In particular, $c=c(\gamma)\rightarrow\frac{M}{n}$ and the first term in (50) vanishes. Therefore, the upper bound converges to a uniform bound on the linear empirical risk.
Corollary D.5.
Under the same assumptions in Corollary D.5, the upper bound on the tilted generalization error in Theorem D.3 has a convergence rate of $\max\big{(}O(1/\sqrt{n}),O(n^{-\beta})\big{)}$ as $n\rightarrow\infty$.
@@
Theorem D.6 (Uniform Lower Bound).
Under the assumptions of Theorem D.3 for a fixed $h\in\mathcal{H}$ we have
@@
Remark D.7.
For $\frac{\log(n)}{2M}<|\gamma|<\frac{\log(n+1)}{M}$, the lower bound in Theorem D.6 for a negative $\gamma$ can be tighter than the lower bound on the tilted generalization error, (53), for a positive $\gamma$, due to the term $\frac{\exp(-2\gamma M)}{n}-1$ in Theorem D.6.
Remark D.8.
The terms $\frac{(1/n-\exp(-2\gamma M))}{\gamma}$ in Theorem D.3 can cause the upper bounds in Theorem D.3 to be tighter for $\gamma>0$ than for $\gamma<0$. Furthermore, in comparison with the uniform upper bound on the linear generalization error, (15), we obtain a tilted generalization error upper bound, (50), which can be tighter, for sufficiently large value of samples $n$ and small values of $\gamma$.
D.2 Information-theoretic bounds: details for bounded loss
@@
The proof follows directly from applying Lemma C.15 to the $\log(.)$ function and then applying Lemma C.14. ∎
@@
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\leq\frac{\exp(-\gamma M)}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\leq\frac{-\exp(-\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(-\gamma M)}{2\gamma}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\exp(-\gamma M)\gamma}{2}(1-1/n)\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}\end{split}$ (56)
@@
$\begin{split}&\mathrm{R}(H,P_{H}\otimes\mu)-\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})\\ &=\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i})))]-\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{\gamma}\log(\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i})))]\\ &\geq\frac{1}{\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Big{[}\exp(\frac{\gamma}{n}\sum_{i=1}^{n}\ell(H,Z_{i}))-\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))\Big{]}\\ &\geq\frac{-\exp(\gamma M)}{2\gamma}\mathbb{E}_{P_{H}\otimes\mu^{\otimes n}}\Bigg{[}\Big{(}\frac{1}{n}\sum_{i=1}^{n}\gamma^{2}\ell(H,Z_{i})^{2}\Big{)}-\Big{(}\frac{1}{n^{2}}\big{(}\sum_{i=1}^{n}\gamma\ell(H,Z_{i})\big{)}^{2}\Big{)}\Bigg{]}\\ &=\frac{-\exp(\gamma M)}{2\gamma}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\gamma\ell(H,\tilde{Z}))\big{]}\\ &=\frac{-\gamma\exp(\gamma M)}{2}\big{(}1-\frac{1}{n}\big{)}\mathbb{E}_{P_{H}}\big{[}\mathrm{Var}_{\tilde{Z}\sim\mu}(\ell(H,\tilde{Z}))\big{]}.\end{split}$ (57)
@@
Corollary 3.15 (Restated).
Under the same assumptions in Theorem 3.13, the absolute value of the expected titled generalization error satisfies $\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|}\sqrt{\frac{I(H;S)}{2n}}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
Proof.
We can derive the upper bound on absolute value of the expected tilted generalization error by combining Theorem 3.13 and Theorem 3.14. ∎
D.2.1 Individual Sample Bound Discussion
We can apply previous information-theoretic bounding techniques (e.g., (Bu et al., 2020), (Asadi et al., 2018), and (Steinke and Zakynthinou, 2020)), exploiting the Lipschitz property of the logarithm function (or Lemma C.10) over a bounded support. For example, to derive an upper bound based on the individual sample (Bu et al., 2020), using the approach for Proposition 3.12 and Lemma C.10, we have for $\gamma>0$, we have
$\begin{split}&\overline{\mathrm{R}}_{\gamma}(H,P_{H}\otimes\mu^{\otimes n})-\overline{\mathrm{R}}_{\gamma}(H,P_{H,S})\\ &\leq\frac{1}{\gamma}\Big{(}\operatorname{E}_{P_{H}\otimes\mu^{\otimes n}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]-\operatorname{E}_{P_{H,S}}[\frac{1}{n}\sum_{i=1}^{n}\exp(\gamma\ell(H,Z_{i}))]\Big{)}\\ &\leq\frac{\big{(}\exp(\gamma M)-1\big{)}}{\gamma}\sum_{i=1}^{n}\frac{1}{n}\sqrt{\frac{I(H;Z_{i})}{2}},\end{split}$ (58)
where for the first equality, we applied the Lemma C.10 and for the second inequality, we use that $1\leq\exp(\gamma\ell(H,Z_{i}))\leq\exp(\gamma M)$ for $\gamma>0$ and the approach in (Bu et al., 2020) for bounding via individual samples. A similar approach also can be applied to $\gamma<0$. Therefore, the following upper bound holds on the absolute value of the expected generalization error,
$\begin{split}|\overline{\mathrm{gen}}_{\gamma}(H,S)|&\leq\frac{(\exp(|\gamma|M)-1)}{|\gamma|n}\sum_{i=1}^{n}\sqrt{\frac{I(H;Z_{i})}{2n}}+\frac{|\gamma|M^{2}\exp(|\gamma|M)}{8}\Big{(}1-\frac{1}{n}\Big{)}.\end{split}$
D.2.2 Another Approach for an Information-theoretic Bound
@@
Using Proposition D.9 and Lemma C.3, we can derive the following upper bound on the expected generalization error.
Theorem D.10.
@@
Using Proposition D.9, we have
@@
Appendix E Proofs and Details of Section 4
E.1 Uniform bounds: details for unbounded loss
@@
Given Assumption 4.1, for any fixed $h\in\mathcal{H}$ with probability at least $(1-\delta)$, then the following upper bound holds on the tilted generalization error for $\gamma<0$, $\begin{split}\mathrm{gen}_{\gamma}(h,S)&\leq 2\kappa_{u}\exp(-\gamma\kappa_{u})\sqrt{\frac{\log(2/\delta)}{n}}-\frac{4\exp(-\gamma\kappa_{u})\log(2/\delta)}{3n\gamma}-\frac{\gamma}{2}\kappa_{u}^{2}.\end{split}$ (67)
@@
Using Bernstein’s inequality, Lemma C.7, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$ and considering $10<X_{i}<$ 1, we have
@@
$\begin{split}&\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\frac{1}{\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]}\frac{4\log(2/\delta)}{3n}\\ \leq&\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}\\ &+\exp(-\gamma\kappa_{u})\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma\kappa_{u})\frac{4\log(2/\delta)}{3n}.\end{split}$
@@
$\begin{split}&\frac{1}{\gamma}\log\Big{(}\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))]\Big{)}-\frac{1}{\gamma}\log\Big{(}\frac{1}{n}\sum_{i=1}^{n}\exp{\gamma\ell(h,Z_{i})}\Big{)}\\ &\leq\frac{\exp(-\gamma\kappa_{u})}{|\gamma|}\sqrt{\frac{4\mathrm{Var}(\exp(\gamma\ell(h,\tilde{Z})))\log(2/\delta)}{n}}+\exp(-\gamma\kappa_{u})\frac{4\log(2/\delta)}{3n|\gamma|}.\end{split}$
Using Lemma C.2, completes the proof. ∎
@@
Given Assumption 4.1, there exists a $\zeta\in(0,1)$ such that for $n\geq\frac{(4\gamma^{2}\kappa_{u}^{2}+8/3\zeta)\log(2/\delta)}{\zeta^{2}\exp(2\gamma\kappa_{u})}$ and $\gamma<0$, for any fixed $h\in\mathcal{H}$, the following lower bound on the tilted generalization error holds with probability at least $(1-\delta)$; $\mathrm{gen}_{\gamma}(h,S)\geq-\frac{2\kappa_{u}\exp(-\gamma\kappa_{u})}{(1-\zeta)}\sqrt{\frac{\log(2/\delta)}{n}}+\frac{4\exp(-\gamma\kappa_{u})(\log(2/\delta))}{3n\gamma(1-\zeta)}.$ (69)
@@
First, we apply Lemma C.9 to yield $\mathrm{R}(h,\mu)-\frac{1}{\gamma}\log(\mathbb{E}[\exp(\gamma\ell(h,\tilde{Z}))])\geq 0$. Next we focus on the second line of this display. Bernstein’s inequality, Lemma C.7, for $X_{i}=\exp(\gamma\ell(h,Z_{i}))$, so that $10<X_{i}<$ 1, gives that with probability at least $(1-\delta)$,